Equation X^3 + 21X^2 + p = 0$$ has three different real solutions that form an arithmetic progression. Find the largest value of parameter p.


1
Expert's answer
2021-09-22T05:21:43-0400

Solution:

"x^3+21x^2+p=0"

It can be written as "(x-a)(x-b)(x-c)=0"

Here, roots are "a,b,c".

Then, "x^3+21x^2+p=(x-a)(x-b)(x-c)"

"\Rightarrow x^3+21x^2+p=(x-a)(x^2-x(b+c)+bc)
\\ \Rightarrow x^3+21x^2+p=x^3-ax^2-x^2(b+c)+ax(b+c)+xbc-abc
\\ \Rightarrow x^3+21x^2+p=x^3+x^2(-a-b-c)+x(ab+ac+bc)-abc"

On comparing both sides,

"-a-b-c=21;\ ab+ac+bc=0;-abc=p
\\\Rightarrow a+b+c=-21\ ...(i);\ ab+ac+bc=0;abc=-p"

It is given that the roots are in AP.

"a+c=2b\ ...(ii)"

using (ii) in (i),

"3b=-21
\\ \Rightarrow b=-7"

Put this value in (ii),

"a+c=-14"

Also, "abc=-p"

"\Rightarrow ac(-7)=-p" [Using b=-7]

"\Rightarrow ac=\dfrac p7"

Also, "ab+ac+bc=0"

"\Rightarrow -7a+ac-7c=0" [Using b=-7]

"\Rightarrow 7(a+c)=ac
\\\Rightarrow 7(-14)=\dfrac p7
\\ \Rightarrow p=-686"